# Knowledge Graph KGE Knowledge Graph Embedding

### 06 Mar 2026 v1

@@
KG
Knowledge Graph
KGE
Knowledge Graph Embedding
FOL
First-Order Logic
AI
Artificial Intelligence
EPFO
Existential Positive First-order
DAG
Directed Acyclic Graph
DNF
Disjunctive Normal Form
AI
Artificial Intelligence
CL
Concept Learning
CEL
Class Expression Learning
CWA
Close World Assumption
CSV
Comma-Separated Values
CNN
Convolutional Neural Network
DL
Description Logic
DNN
Deep Neural Network
FOL
First-Order Logic
GD
Gradient Descent
KB
Knowledge Base
KG
Knowledge Graph
KGE
Knowledge Graph Embedding
SW
Semantic Web
OWA
Open World Assumption
SGD
Stochastic Gradient Descent
KGC
Knowledge Graph Completion
MAP
Maximum a Posteriori Probability Estimation
MRR
Mean Reciprocal Rank
ML
Machine Learning
MLE
Maximum Likelihood Estimation
MDP
Maximum Likelihood Estimation
MTL
Multi-task Learning
MDP
Markov Decision Process
NN
Neural Network
KP
Kronecker Product
KD
Kronecker Decomposition
KB
Knowledge Base
KG
Knowledge Graph
KGE
Knowledge Graph Embedding
ILP
Inductive Logic Programming
RL
Reinforcement Learning
RNN
Recurrent Neural Network
RDF
Resource Description Framework
SWA
Stochastic Weight Averaging
OWL
Web Ontology Language
OWA
Open World Assumption
WWW
World Wide Web
11institutetext: Heinz Nixdorf Institute, Furstenalle 11, 33102 Paderborn, Germany 22institutetext: Paderborn University, Warburgerstr. 100, 33098 Paderborn, Germany
Semantics-Aware Caching for Concept Learning
Louis Mozart Kamdem Teyou    Caglar Demir &Axel-Cyrille Ngonga Ngomo
[…]
2Paderborn University
[…]
   cdmir    axel.ngonga}@upb.de    Louis Mozart Kamdem Teyou    Caglar Demir    Axel-Cyrille Ngonga Ngomo louis888    cdmir    axel.ngonga}@upb.de
@@ Abstract
Concept learning is a form of supervised machine learning that operates on knowledge bases in description logics. State-of-the-art concept learners often rely on an iterative search through a countably infinite concept space. In each iteration, they retrieve instances of candidate solutions to select the best concept for the next iteration. While simple learning problems might require a few dozen instance retrieval calls to find a fitting solution, complex learning problems might necessitate thousands of calls. We alleviate the resulting runtime challenge by presenting a semantics-aware caching approach. Our cache is essentially a subsumption-aware map that links concepts to a set of instances via crisp set operations. Our experiments on 5 datasets with 4 symbolic reasoners, a neuro-symbolic reasoner, and 5 popular pagination policies demonstrate that our cache can reduce the runtime of concept retrieval and concept learning by an order of magnitude while being effective for both symbolic and neuro-symbolic reasoners.
@@ 1 Introduction
Knowledge bases are now first-class citizens of the Web, with roughly 50% of websites hosting assertions from RDF knowledge bases. [^1] https://webdatacommons.org/structureddata/ Given the advent of machine learning across disciplines and on corpora of increasing size, it stands to reason that the over $10^{11}$ assertions available on the Web need to be made amenable to machine learning. Embedding approaches and geometric deep learning are used to map the constituents of knowledge bases to vectorial representations [32]. However, these approaches do not make explicit of the semantics of RDF knowledge bases. In contrast, concept learning (also called class expression learning, short: CEL) [10, 25, 14] is designed to exploit the explicit semantics of description logics to derive ante-hoc explainable models.
The goal of CEL is, given a set of positive and negative examples, to describe the set of positive examples and not the set of negative examples. This task is equivalent to learn a class expression that describes individuals provided as positive examples [20, 19]. This problem is inherently challenging, as most CEL algorithms operate over a quasi-ordered infinite hypothesis space defined by the expressivity of the underlying description logic (short: DL). The majority of state-of-the-art CEL methods adopt a top-down refinement strategy, which initiates the search from the most general concept (e.g., $\top$) and incrementally specializes it through a sequence of refinement steps [20]. This iterative process continues until a hypothesis that best approximates the target concept in terms of covering positive examples while avoiding negatives is reached.
@@
A central bottleneck in CEL is the repeated use of DL reasoners to evaluate candidate class expressions. Each candidate must be passed to a reasoner to retrieve its instances, an operation that is both computationally expensive and invoked thousands of times during the search. Accelerating this evaluation step, therefore, has a direct and substantial impact on the overall runtime of CEL algorithms. This work targets this bottleneck by accelerating the reasoner performance as a foundational component of efficient CEL.
We introduce a semantic caching mechanism designed to accelerate instance retrieval in Description Logic reasoners. Our contribution can be summarized as follows: regardless of the underlying reasoner our approach can substantially improve performance by up to $80\%$ for slower reasoners and by up to $20\%$ for faster ones, provided that sufficient cache capacity [^2] In our experiments, cache capacity refers to the number of distinct class expressions and their retrieved instances that can be stored in the cache dictionary, rather than physical memory usage. is available.
Traditional approaches, such as ELK [15], demonstrate how materialization-based techniques can drastically improve reasoning efficiency by precomputing subsumptions and class memberships. While our work shares with ELK the overarching goal of scalability, the two approaches are fundamentally different. ELK is itself a highly optimized reasoner restricted to the $\mathcal{EL}^{++}$ description logic. In contrast, our method is not a new reasoner but rather a complementary caching layer that accelerates existing reasoning tasks. For this reason, a direct experimental comparison is not meaningful. Instead, our approach leverages the semantics of $\mathcal{ALC}$ to systematically compose the set of instances of complex concepts. By reusing previously computed results through caching, our method avoids redundant reasoning calls and significantly reduces the computational cost of instance retrieval of a given reasoner.
We demonstrate that this mechanism can be integrated with little effort into existing reasoning pipelines by evaluating the performance of 4 symbolic and 1 neuro-symbolic reasoners with and without our cache within the setting of concept retrieval and concept learning. Our results suggest that our caching approach can also reduce concept learning runtimes by up to three orders of magnitude. Our codes, datasets, and experimental setup can be found in the supplementary material  [^3] https://anonymous.4open.science/r/Cache-algorithm-for-concept-learners-F772/.
@@ 2.1 Concept learning algorithms
EvoLearner [12], a more recent development, takes an evolutionary approach to CEL. Instead of relying only on refinement operators, EvoLearner initializes its population of candidate expressions through random walks on the knowledge graph, which are then converted into description logic concepts represented as abstract syntax trees. These candidate concepts are refined further using mutation and crossover operations, enabling EvoLearner to outperform traditional refinement operator-based methods like CELOE and OCEL in both efficiency and accuracy. Another example is ECII [26], a search-based algorithm that does not employ refinement operators, invoking a reasoner only once during execution. This design choice allows ECII to bypass the iterative computational overhead which is common in traditional methods.
@@
Another promising approach is Neuro-Symbolic Class Expression Learning (DRILL) [8], which addresses the limitations of heuristic-based methods like CELOE and OCEL. Traditional methods rely heavily on myopic heuristics and fixed rules for redundancy elimination and expression simplification, which can lead to suboptimal runtimes and memory inefficiencies. DRILL leverages deep Q-learning to dynamically adapt the search process, reducing runtimes and memory requirements while enhancing scalability. This reinforcement learning paradigm overcomes the drawbacks of fixed heuristic rules, providing a more flexible and efficient framework for CEL.
@@
Despite recent progress, scaling CEL methods remains a major challenge due to persistent memory and runtime constraints [8, 17]. Caching mechanisms offer a promising solution by optimizing instance retrieval—the primary computational bottleneck in CEL. The following subsection explores caching algorithms and eviction strategies aimed at improving the scalability and efficiency of CEL systems.
@@ 2.2 Cache algorithms with replacement policies
In our experiments, we aimed to evaluate the impact of policies based on recency and order of entry on the performance of reasoners, and we observed that policies based on frequency yielded similar results to recency based—specifically, MRU behaved similarly to MFU, and LRU to LFU. Additionally, we investigated the Random Placement (RP) strategy, which involves randomly evicting an object from the cache to make space for a new one. While modern approaches, such as those leveraging machine learning, deep learning, and reinforcement learning [27, 33], have been explored to enhance cache replacement policies, these methods fall outside the scope of our study.
@@ 2.3 Metrics
Our work builds on these foundations by introducing a simple yet powerful caching mechanism tailored specifically for semantic reasoning tasks. By focusing on the subsumption relationships between concepts, we provide a novel perspective on integrating caching into CEL pipelines. The next section delves deeper into our proposed algorithms with a focus on how they can be used to accelerate instance retrieval in CEL.
@@ 3.2 Class Expression Learning
Let $N_{I}$ be a set of individuals. Given a KB $\mathcal{K}$, a description logic $\mathcal{L}$ (e.g., $\mathcal{ALC}$, $\mathcal{SHOIN}$, $\mathcal{SROIQ}$), and sets $E^{+}\subseteq N_{I}$ of positive and $E^{-}\subseteq N_{I}$ of negative examples, the goal of CEL is to construct a concept $H$ in $\mathcal{L}$ such that $\forall e^{+}\in E^{+}:\mathcal{K}\models H(e^{+})$ and $\forall e^{-}\in E^{-}:\mathcal{K}\not\models H(e^{-})$ [17, 8, 10]. A large proportion of the approaches to CEL [20, 19, 10, 25, 14] rely on iterative search through the infinite space $\mathcal{C}(\mathcal{L})$ of all concepts that can be built in $\mathcal{L}$. In each iteration, they compute candidate concepts $C$ and their instances $C$ by issuing an instance retrieval query to a reasoner. As pointed out by [14], this instance retrieval step is the main runtime bottleneck for concept learners. In particular, concept learners may need to retrieve the instances of thousands of concepts during each iteration, further exacerbating the computational load due to instance retrieval.
Notation
@@ 4 Approach
While our approach is inherently adaptable to more expressive DL, this work specifically concentrates on the $\mathcal{ALC}$ DL. The syntax and semantics of $\mathcal{ALC}$ are presented in Table 1.
@@ 4.1 Overview
Our approach is based on the definition of subsumption: Let $C$ and $D$ be two concepts. If $C\sqsubseteq D$, then $Ret(C)\subseteq Ret(D)$. While a naïve instance retrieval would check whether $C(a)$ holds for all $a\in N_{I}$ to compute $Ret(C)$, we are guaranteed to complete the same computation by checking solely $|Ret(D)|$ individuals. Our intuition is that if $|Ret(D)|<<|N_{I}|$, then the computation of $Ret(C)$ should be more time-efficient using our cache.
The formulation of our intuition elucidates the main challenge behind our approach, i.e., finding the most specific $D$ with $C\sqsubseteq D$. Determining whether $C\sqsubseteq D$ is known to be NP-hard in $\mathcal{ALC}$ [7] hence, we adopt the following set of heuristics:
  • $C\equiv D\sqcap E\rightarrow C\sqsubseteq D$
  • $C\equiv\exists r.C\wedge D\equiv\exists r.\top\rightarrow C\sqsubseteq D$
  • $C\sqsubseteq D_{1}\wedge\ldots\wedge C\sqsubseteq D_{n}\rightarrow C\sqsubseteq\bigsqcap\limits_{i=1}^{n}D_{i}$
The correctness of these heuristics is a direct consequence of the definition of concept subsumption. Based on these heuristics, we can now describe how our cache, which we denote $\mathcal{A}$, works. Given a concept $C$, we begin by searching for all $D_{i}$ in $\mathcal{A}$’s key set with $C\sqsubseteq D_{i}$ using the heuristics aforementioned. Then, we retrieve $\bigcap\limits_{i}D_{i}$ and check whether $C(a)$ holds for every individual $a\in\bigcap\limits_{i}D_{i}$. In particular, if we can deduce $C\equiv D$ for any $D$ in $\mathcal{A}$’s key set, then we simply fetch $Ret(D)$ from $\mathcal{A}$. The details of our approach are described in Algorithm 3 and the subsequent sections.
Although the semantic cache algorithm can operate without an initialization phase, we argue that performing this step is almost always worthwhile. When time permits, especially for datasets that are frequently used—it is beneficial to perform this step to improve effectiveness. We add the following precomputed instance retrieval results to $\mathcal{A}$:
@@ 4.2 Fetching from Cache and Space Management
The overall space management of the cache is handled by Algorithm 2 while the cache fetch is governed by Algorithm 3, which allows for the retrieval of instances of concepts. The algorithm begins by checking whether the queried concept $C$ is a named concept. In this case, its instances are already precomputed and can be returned immediately without further reasoning. If $C$ is not a named concept, the algorithm considers trivial cases (lines 2–5).
If the concept is neither named nor trivial, the algorithm proceeds recursively by decomposing the concept according to its semantics. For negations (line 6), the instances of the sub-concept are retrieved, and the complement with respect to N_IN\_I is taken. For conjunctions (line 8), the intersection of the instances of $D$ and $E$ is computed, while for disjunctions (line 8), their union is returned. Existential restrictions (lines 12–23) are handled by first computing the instances of the filler concept $D$ and then iterating over all individuals $a\in N_{I}$ to check, for each candidate $b\in$ instances $(D)$, whether the reasoner confirms that $r(a,b)$ holds. If such a witness is found, $a$ is added to the result set. Universal restrictions (line 24) are reduced to a negated existential restriction, i.e., $\neg(\exists r.\neg D)$, which is then handled by the previous case.
Once the recursive decomposition has been exhausted, the algorithm checks whether the result for $C$ is already stored in the cache. If it is not, the reasoner is invoked to compute the instances of $C$, and the result is stored in the cache as a new entry (lines 28-30). This ensures that subsequent queries for the same concept can be answered without repeating the reasoning process. Finally, the result is returned from the cache (line 32).
Cache space management is handled by the Algorithm 2. When a new result $R$ is to be stored, the algorithm first checks whether the cache can accommodate it (line 1). If the maximum size is exceeded, a purge operation is triggered to free space (line 2). Afterwards, the pair $(C,R)$ is added to the cache (line 4).

### 12 Mar 2026 v2

# Semantics-Aware Caching for Concept Learning

@@ Abstract
Concept learning is a form of supervised machine learning that operates on knowledge bases in description logics. State-of-the-art concept learners often rely on an iterative search through a countably infinite concept space. In each iteration, they retrieve instances of candidate solutions to select the best concept for the next iteration. While simple learning problems might require a few dozen concept retrieval calls to find a suitable solution, complex learning problems might require thousands of calls, making concept retrieval a major runtime bottleneck. We alleviate the resulting runtime challenge by presenting a semantics-aware caching approach. Our cache is essentially a subsumption-aware map that links concepts to a set of instances via crisp set operations. We evaluate our approach on four datasets using four symbolic reasoners, one neuro-symbolic reasoner, and five cache eviction policies. The results show that our cache reduces concept retrieval and concept learning runtimes by up to an order of magnitude while remaining effective across both symbolic and neuro-symbolic reasoning reasoners.
@@ 1 Introduction
Knowledge bases have become first-class citizens of the Web, with roughly 50% of websites containing structured data expressed in RDF knowledge bases. [^1] https://webdatacommons.org/structureddata/ Given the advent of machine learning across disciplines and on corpora of increasing size, it stands to reason that the over $10^{11}$ assertions available on the Web need to be made amenable to machine learning. Embedding approaches and geometric deep learning are used to map the constituents of knowledge bases to vectorial representations [32].
While effective for many predictive tasks, these methods typically approximate the relational structure of knowledge bases and do not explicitly preserve the logical semantics of RDF and description logics. In contrast, concept learning (also called class expression learning, short: CEL) [10, 25, 14] operates directly on symbolic representations and produces interpretable models in the form of description logic class expressions.
The goal of CEL is to learn a class expression that describes a set of positive examples while excluding negative ones [20, 19]. In practice, this requires searching in an infinite, quasi-ordered hypothesis space defined by the expressivity of the underlying description logic (short: DL). Most state-of-the-art CEL algorithms use a top-down refinement operator that starts from the most general concept (e.g., $\top$) and iteratively specializes it through a sequence of refinement steps [20]. This iterative process continues until a hypothesis that best approximates the target concept in terms of covering positive examples while avoiding negatives is reached.
@@
A central computational bottleneck of CEL is the repeated use of description logic reasoners to evaluate candidate class expressions. Each candidate must be submitted to a reasoner to retrieve its instances, an operation that can be computationally expensive. During the search process, hundreds or even thousands of candidate expressions may be evaluated, resulting in many repeated instance-retrieval calls. Consequently, the efficiency of concept retrieval has a direct impact on the overall runtime of concept learning algorithms.
Despite the importance of this operation, we could not identify prior work that systematically investigates caching mechanisms for accelerating concept retrieval. In this work, we address this bottleneck by introducing a semantics-aware caching mechanism designed to accelerate concept retrieval in description logic reasoners. Our cache can be viewed as a subsumption-aware map that associates class expressions with their corresponding sets of instances. By exploiting the semantics of $\mathcal{ALC}$, the cache can derive the instances of new concepts from previously computed ones through simple set operations, thereby avoiding redundant reasoning calls.
Traditional approaches such as ELK [15] demonstrate how materialization-based techniques can significantly improve reasoning efficiency by precomputing subsumptions and class memberships. Although our work shares the general goal of improving scalability, the two approaches differ fundamentally. ELK is an optimized reasoner designed specifically for the $\mathcal{EL}^{++}$ description logic. In contrast, our approach does not introduce a new reasoner. Instead, it provides a complementary caching layer that accelerates concept retrieval for existing reasoners without modifying their internal algorithms. For this reason, a direct experimental comparison with ELK is not meaningful.
We evaluate the proposed caching mechanism on four datasets using four symbolic reasoners and one neuro-symbolic reasoner. Experiments are conducted for both concept retrieval and full concept-learning pipelines. Our results show that the proposed cache can substantially improve performance, reducing concept retrieval time by up to $80\%$ for slower reasoners and by up to $20\%$ for faster ones when sufficient cache capacity is available. When integrated into concept learning systems, the cache can reduce overall learning runtimes by up to three orders of magnitude.
Our code, datasets, and experimental setup are available in the supplementary material. [^2] https://github.com/Louis-Mozart/Cache-algorithm-for-concept-learners
@@ 2.1 Concept Learning Algorithms
EvoLearner [12], a more recent development, takes an evolutionary approach to CEL. Instead of relying only on refinement operators, EvoLearner initializes its population of candidate expressions through random walks on the knowledge graph, which are then converted into description logic concepts represented as abstract syntax trees. These candidate concepts are refined further using mutation and crossover operations, enabling EvoLearner to outperform traditional refinement operator-based methods like CELOE and OCEL in both efficiency and accuracy. Another example is ECII [26], a search-based algorithm that does not employ refinement operators, invoking a reasoner only once during execution. This design choice allows ECII to bypass the iterative computational overhead that is common in traditional methods.
@@
Another promising approach is Neuro-Symbolic Class Expression Learning (DRILL) [8], which addresses the limitations of heuristic-based methods like CELOE and OCEL. Traditional methods rely heavily on myopic heuristics and fixed rules for redundancy elimination and expression simplification, which can lead to suboptimal runtimes and memory inefficiencies. DRILL uses deep Q-learning to dynamically adapt the search process, reducing runtimes and memory requirements while enhancing scalability. This reinforcement learning paradigm overcomes the drawbacks of fixed heuristic rules, providing a more flexible and efficient framework for CEL.
@@
Despite recent progress, scaling CEL methods remains a major challenge due to persistent memory and runtime constraints [8, 17]. Caching mechanisms offer a promising solution by optimizing concept retrieval—the primary computational bottleneck in CEL.
@@ 2.3 Metrics
Our work builds on these foundations by introducing a simple caching mechanism tailored specifically for semantic reasoning tasks. By focusing on the subsumption relationships between concepts, we provide a novel perspective on integrating caching into CEL pipelines. The next section delves deeper into our proposed algorithms with a focus on how they can be used to accelerate concept retrieval in CEL.
@@ 3.2 Class Expression Learning
Let $N_{I}$ denote the set of individuals in a knowledge base $\mathcal{K}$. Given a description logic $\mathcal{L}$ (e.g., $\mathcal{ALC}$, $\mathcal{SHOIN}$, or $\mathcal{SROIQ}$) and two sets of labeled examples $E^{+}\subseteq N_{I}$ (positive examples) and $E^{-}\subseteq N_{I}$ (negative examples), the goal of CEL is to construct a concept $H\in\mathcal{C}(\mathcal{L})$ such that
$\forall e^{+}\in E^{+}:\mathcal{K}\models H(e^{+})\quad\text{and}\quad\forall e^{-}\in E^{-}:\mathcal{K}\not\models H(e^{-}),$
where $\mathcal{C}(\mathcal{L})$ denotes the set of all concepts that can be expressed in $\mathcal{L}$ [17, 8, 10]. In practice, perfect separation is often impossible, and learners therefore aim to identify concepts that optimize a quality measure such as accuracy, precision, recall, or F1-score.
Most CEL algorithms explore the concept space $\mathcal{C}(\mathcal{L})$ through heuristic search guided by refinement operators [20, 19]. A common approach is top-down search, where the process starts from the most general concept (e.g., $\top$) and incrementally specializes it by applying refinement rules that introduce logical constructors such as conjunctions, disjunctions, or existential restrictions. Other approaches explore the space bottom-up or combine symbolic search with neural guidance [10, 25, 17, 12].
Regardless of the search strategy, evaluating candidate concepts is a central step in CEL. For each generated concept $C$, the learner must determine the set of individuals $C^{I}$ that satisfy the concept in $\mathcal{K}$. This operation corresponds to concept retrieval and typically requires invoking a description logic reasoner. The retrieved instances are then compared with $E^{+}$ and $E^{-}$ to compute the quality of the candidate concept. Prior work has identified concept retrieval as one of the main computational bottlenecks in CEL pipelines [14]. This observation motivates the development of mechanisms that can reuse previously computed reasoning results to reduce redundant calls to the reasoner.
Notation:
@@ 4 Approach
Our goal is to accelerate concept retrieval in description logic reasoners by introducing a semantics-aware cache. While the approach can in principle be extended to more expressive logics, we focus on $\mathcal{ALC}$ in this work because it provides the core constructors required by many CEL systems while keeping the exposition clear. The syntax and semantics of $\mathcal{ALC}$ concepts are summarized in Table 1.
@@ 4.1 Overview
Our approach exploits a fundamental property of description logics: if a concept $C$ is subsumed by another concept $D$, then every instance of $C$ must also be an instance of $D$, i.e.,
$C\sqsubseteq D\Rightarrow Ret(C)\subseteq Ret(D).$
A naive concept retrieval procedure evaluates $C(a)$ for every individual $a\in N_{I}$ to compute $Ret(C)$. However, if we already know the set $Ret(D)$ for a concept $D$ with $C\sqsubseteq D$, we only need to verify the membership of individuals in $Ret(D)$ to determine $Ret(C)$. When $|Ret(D)|\ll|N_{I}|$, this can significantly reduce the number of reasoning calls.
This observation motivates our caching strategy. Instead of recomputing concept retrieval queries from scratch, the cache stores previously evaluated concepts together with their instance sets. When a new query concept $C$ is issued, the cache attempts to identify previously stored concepts that subsume $C$ and uses their instance sets to restrict the search space.
The main challenge lies in identifying such concepts efficiently. Determining whether $C\sqsubseteq D$ is NP-hard in $\mathcal{ALC}$ [7]. Performing general subsumption checks during cache lookup would therefore defeat the purpose of the cache. Instead, we employ a small set of sound heuristics that detect subsumption relationships through simple syntactic patterns. These heuristics operate only on concepts already stored in the cache and avoid invoking a reasoner.
4.2 Subsumption Heuristics and Correctness
The heuristics used by our cache are designed to identify subsumption relations through simple structural patterns in concept expressions. Importantly, whenever a heuristic concludes that $C\sqsubseteq D$, the subsumption relation is guaranteed to hold under the standard semantics of description logics.
The heuristics used in our implementation are the following:
  • If $C\equiv D\sqcap E$, then $C\sqsubseteq D$.
  • If $C\equiv\exists r.C^{\prime}$ and $D\equiv\exists r.\top$, then $C\sqsubseteq D$.
  • If $C\sqsubseteq D_{1},\ldots,C\sqsubseteq D_{n}$, then $C\sqsubseteq\bigsqcap_{i=1}^{n}D_{i}.$
These rules follow directly from the semantics of $\mathcal{ALC}$. For instance, the first rule holds because the interpretation of a conjunction is defined as the intersection of the interpretations of its conjuncts. Similarly, the existential rule follows from the fact that $\exists r.C^{\prime}$ always implies $\exists r.\top$.
During cache lookup, these heuristics are applied to concepts already present in the cache. The cache is implemented as a hash map whose keys correspond to the abstract syntax trees of concept expressions, see Algorithm 3. The heuristics are evaluated over these keys to identify potential superconcepts of the queried concept.
Importantly, the system does not perform general subsumption reasoning. All checks are local and syntactic, ensuring that cache lookup remains computationally lightweight while preserving correctness.
Once candidate superconcepts $D_{i}$ have been identified in the cache, their instance sets are intersected to obtain a reduced candidate set:
$S=\bigcap_{i}Ret(D_{i}).$
Instead of evaluating $C(a)$ for all $a\in N_{I}$, the reasoner only needs to check individuals in $S$. If an exact match $C\equiv D$ is found among the cache keys, the cache can directly return $Ret(D)$ without invoking the reasoner.
This recursive decomposition, combined with subsumption-based pruning, yields a constructive correctness argument: every individual returned by the cache satisfies the semantics of the queried concept, while individuals outside the candidate set are guaranteed not to belong to the result.
4.3 Cache Initialization
Although the semantic cache can operate without an initialization phase, pre-populating the cache with frequently used concepts significantly improves its effectiveness. In particular, CEL systems often generate candidate expressions built from atomic concepts and simple role restrictions. Precomputing these instance sets, therefore, increases the probability that subsequent queries can reuse cached results.
When initialization is performed, we add the following precomputed concept retrieval results to the cache:
@@ 4.4 Fetching from Cache and Space Management
Algorithms 3 and 2 describe how concept retrieval queries are answered using the cache and how cache entries are managed. The procedure fetchInstances first checks whether the queried concept $C$ is already stored in the cache. For named concepts, the corresponding instance sets are typically precomputed during initialization and can therefore be returned immediately.
If the concept is not directly available, the algorithm evaluates it by recursively decomposing the expression according to the semantics of $\mathcal{ALC}$. Negations are handled by computing the complement of the retrieved instances, conjunctions and disjunctions correspond to set intersection and union, respectively, and existential restrictions are evaluated by checking for role successors that satisfy the filler concept. Universal restrictions are reduced to negated existential restrictions, i.e., $\forall r.C\equiv\neg(\exists r.\neg C)$.
If the result for $C$ is not already stored in the cache, the reasoner is invoked to compute its instance set, and the resulting pair $(C,Ret(C))$ is inserted into the cache. This ensures that subsequent queries involving the same concept can reuse previously computed results.
Cache capacity is controlled by Algorithm 2. When inserting a new entry, the algorithm first checks whether sufficient space is available. If the maximum capacity would be exceeded, a purge operation removes entries according to the selected eviction policy. The new concept and its instance set are then stored in the cache.